. What will be the unit digit of the squares of the following numbers? i. 81 ii. 272 iii. 799  
gyklutld7tdltld

Grade:12th pass

1 Answers

Chetan Jaiswal
2 years ago
Hello! 
 
  1. 81 : (1*1=1), so the units digit of the square of 81 will be 1
  2. 272 : (2*2=4), so the units digit of the square of 272 will be 4
  3. 799 : (9*9=81), so the units digit of the square of 799 will be 1
Thanks!